# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v2

@@ 2.1. Cut-and-paste procedure: oriented setting
We finish this section with the following observation regarding lifts of diffeomorphisms to coverings, whose proof is left to the reader.
@@
be a diffeomorphism between two four-manifolds $M_{1}$ and $M_{2}$ with fundamental group $\pi_{1}(M_{i})=\mathbb{Z}/2p$ and let $\widetilde{M}_{i}$ be the degree-two covering of $M_{i}$ for $i=1,2$. The diffeomorphism (2.8) lifts to a diffeomorphism
@@ 3.1. Raw materials for Theorem C
We will employ as a raw material the exotic four-manifolds $\mathcal{E}$ and $\mathcal{R}$ constructed in [29, Theorem 1.2].
@@
    There exists a smooth four-manifold $\mathcal{E}$, which is homeomorphic but not diffeomorphic to ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$.
@@
    There exists a smooth four-manifold $\mathcal{R}$, which is homeomorphic but not diffeomorphic to $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2$.
We refer the reader to [29, Section 6.2] for the precise definition of the four-manifolds $\mathcal{E}$ and $\mathcal{R}$. The four-manifolds of Theorem C are defined by applying the construction described in Section 2.1 to the exotic four-manifold $\mathcal{R}$ in the second item of Theorem 10.
For any integer $p\geq 1$, we define the closed oriented smooth four-manifold
(3.1) $\mathcal{R}_{2p}=(\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}))\cup S_{p}$
@@
(3.2) $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}},$
@@
where $\alpha_{\Sigma_{2}}\subset\Sigma_{2}$ is a simple loop whose homotopy class generates the fundamental group $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2\cong\pi_{1}(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})$. The second clause of Theorem 10 implies that there is a homeomorphism
$(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\alpha_{\Sigma_{2}})\rightarrow\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}).$
Such homeomorphism can be extended to a homeomorphism as in (3.2) by the decomposition of $\Sigma_{2p}$ in (3.3) and Corollary 5.
Remark 12.
Alternatively, the existence of a homeomorphism as in Lemma 11 follows from work of Hambleton-Kreck [21, Theorem C]; cf. [34, Theorem 2.1].
To distinguish the diffeomorphism class of $\mathcal{R}_{2p}$ from the one of the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, we look at the double covering
@@
corresponding to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p=\pi_{1}(\mathcal{R}_{2p})$.
@@
If $p\in\mathbb{Z}_{>0}$ is odd, the degree-two covering of $\mathcal{R}_{2p}$ corresponding to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to the four-manifold
@@
where $\mathcal{E}$ is the exotic four-manifold of the first clause of Theorem 10, $\gamma\subset\mathcal{E}$ is a simple loop and $S_{p}$ is the four-manifold defined in (2.1).
[Proof]This follows from the definition of $\mathcal{R}$ in [29] as the quotient of $\mathcal{E}$ by a fixed-point-free orientation-reversing involution, together with Lemma 8.
@@ 3.2. Diffeomorphism obstruction
In light of the definition of $\mathcal{E}$ in [29, Section 6.2], Proposition 13 implies the existence of a decomposition
@@
where $V=C_{0}\cup_{\phi_{0}\circ\sigma}Z$ is the building block defined in [29, Construction 6.4]. In particular, $C_{0}$ is a contractible four-manifold, $Z$ consists of five $-1$-framed $2$-handles attached to $\partial C_{0}$, $\partial V$ is the result of 00-surgery on $S^{3}$ along the Stevedore knot and $\zeta$ is any of its orientation-reversing fixed-point-free involutions. On the other hand, for every $p\geq 1$, $V_{p}$ is defined to be the four-manifold
@@
where $C_{0}^{p}$ is the four-manifold obtained from $C_{0}$ by removing a neighbourhood of a simple closed curve and gluing back a copy of $S_{p}$. Moreover, we set
$\mathcal{E}^{\prime}_{p}=V^{\prime}\cup_{\zeta}V^{\prime}_{p}$
where $V^{\prime}=C_{0}\cup_{\phi_{0}}Z$ and $V^{\prime}_{p}=C_{0}^{p}\cup_{\phi_{0}}Z$ are made up by the same building blocks as $V$ and $V_{p}$ with different gluing maps. We refer the reader to [29, Construction 6.4] for further details.
@@
    $X_{p}=(({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\mathbin{\#}\Sigma_{p},$
    where $\gamma\subset{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ is a simple loop.
@@
We proceed to provide some technical lemmas that are needed to distinguish the smooth structures of Theorem C. These results are generalizations of [29, Proposition 6.9] and [29, Lemma 6.3] respectively. In particular, we make use of the Ozsváth-Szabó four-manifold invariants defined in [31, Section 2.4].
In the following, we denote by $\Psi_{W,\mathfrak{t}}\in HF^{-}_{\text{red}}(\partial W,\mathfrak{t}|_{\partial W})$ the relative invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(W,\mathfrak{t})$ where $W$ is a compact four-manifold with non-empty boundary, and $\Phi_{X,\mathfrak{s},L}$ is the closed four-manifold invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with respect to a one-dimensional subspace $L\subset H_{2}(X;\mathbb{Q})$. We refer the reader to [29, Section 6.1] for the needed background on such invariants.
@@
For any $p\geq 1$, there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(V_{p},\mathfrak{t}_{p})$ such that $\Psi_{V_{p},\mathfrak{t}_{p}}\neq 0$. Moreover, all such ${\operatorname{Spin}}^{\mathbb{C}}$-structures can be chosen to coincide on $Z\subset V_{p}$. In particular, $V_{p}$ is homeomorphic but not diffeomorphic to $X_{0}(Q_{0})_{p}^{\prime}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for
@@
where $X_{0}(Q_{0})$ is the 00-trace of the Stevedore knot and $\gamma\subset X_{0}(Q_{0})$ is a null-homotopic simple loop.
[Proof]The case $p=1$ is [29, Proposition 6.9]. Since $C_{0}^{p}\setminus B^{4}$ is a rational homology cobordism between $S^{3}$ and $\partial C_{0}^{p}$, the same proof holds for $p>1$ by replacing $C_{0}$ with $C_{0}^{p}$ in the argument. The proof of the following result is a straightforward adaptation of the one of [29, Lemma 6.3].
@@
Let $K$ be a fibered amphichiral knot of genus $g$. Let $W_{1}$ and $W_{2}$ be compact oriented 4-manifolds with positive second Betti number and boundary $S^{3}_{0}(K)$. Suppose there are ${\operatorname{Spin}}^{\mathbb{C}}$ structures $\{(W_{i},\mathfrak{u}_{i}):i=1,2\}$ such that $c_{1}(\mathfrak{u}_{i})$ evaluates to $22g-$ 2 on a generator of $H_{2}(S^{3}_{0}(K))$ and for which $\Psi_{W_{i},\mathfrak{u}_{i}}\neq 0$. Let $X$ be the result of gluing together $W_{1}$ and $W_{2}$ via any choice of orientation-reversing diffeomorphism of $S^{3}_{0}(K)$, and let $L\subset H_{2}(X;\mathbb{Q})$ be the line corresponding to the inclusion
@@
Then there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(X,\mathfrak{t})$ for which $\Phi_{X,\mathfrak{t},L}\neq 0$.
[Proof of Proposition 14] Item (1) follows from the construction and the fact that $\mathcal{E}^{\prime}$ is an exotic ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ by [29, Theorem 6.7]. By Lemma 7, there is a degree-two covering
$X_{p}\cong\mathcal{E}^{\prime}_{p}\xrightarrow{2:1}\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}.$
Let us now prove Item (2). In particular, $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ are homeomorphic by construction, since $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are homeomorphic by [29, Theorem 6.7]. To conclude the proof, we need to show that the degree-two coverings $\mathcal{E}_{p}\rightarrow\mathcal{R}_{2p}$ and $\mathcal{E}^{\prime}_{p}\rightarrow\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ are not diffeomorphic. Consider the decomposition
@@
where $V_{a}$ and $V_{b}^{p}$ are copies of $V$ and $V_{p}$ respectively without the $T^{-}_{Q_{0}}$ component, see [29, Construction 6.4]. As for the case of $\mathcal{E}$, we see that $H_{2}(\mathcal{E}_{p})$ has a basis
@@
  • $(\Xi,\Theta)$ is the basis for $H_{2}(T_{Q_{0}})$ from [29, Remark 3.11]. In particular, $\Xi$ is represented by a capped-off Seifert surface for $Q_{0}$ in $S^{3}(Q_{0})$ and $\Theta$ is represented by a $2$-sphere of self-intersection $-2$ and $\Xi\cdot\Theta=1$.
As a consequence of Lemma 15 and 16, there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(\mathcal{E}_{p},\mathfrak{t})$ on such that $\Phi_{\mathcal{E}_{p},\mathfrak{t},L}\neq 0$, where $L=\text{Span}(\Xi)\subset H_{2}(\mathcal{E}_{p};\mathbb{Q})$. By [29, Lemma 6.2], we have that $\Xi$ cannot be represented by any surface of genus less than $2$. It is possible to show that if there was a diffeomorphism
$\mathcal{E}_{p}\rightarrow X_{p}$
then one could build a spherical representative for $\Xi$ as in the proof of [29, Theorem 6.7], which would lead to a contradiction.
@@
When $p$ is even, the four-manifold $D_{p}$ of Lemma 2 has two boundary components and the double covering $\mathcal{E}_{p}$ of $\mathcal{R}_{2p}$ is
@@
In particular, we have not been able to show that $\mathcal{R}_{2p}$ is exotic by reducing to the study of the Heegaard Floer invariants of $\mathcal{E}_{p}$ as in [29].

### v3

@@ 2.1. Cut-and-paste procedure: oriented setting
We finish this section with the following observation regarding lifts of diffeomorphisms to covers, whose proof is left to the reader.
@@
be a diffeomorphism between two four-manifolds $M_{1}$ and $M_{2}$ with fundamental group $\pi_{1}(M_{i})=\mathbb{Z}/2p$ and let $\widetilde{M}_{i}$ be the degree-two cover of $M_{i}$ for $i=1,2$. The diffeomorphism (2.13) lifts to a diffeomorphism
@@ 3.1. Raw materials for Theorem C
We will employ as a raw material the exotic four-manifolds $\mathcal{E}$ and $\mathcal{R}$ constructed in [32, Theorem 1.2].
@@
    There exists a four-manifold $\mathcal{E}$, which is homeomorphic but not diffeomorphic to ${\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$.
@@
    There exists a four-manifold $\mathcal{R}$, which is homeomorphic but not diffeomorphic to $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2$.
We refer the reader to [32, Section 6.2] for the precise definition of the four-manifolds $\mathcal{E}$ and $\mathcal{R}$. The four-manifolds of Theorem C are defined by applying the construction described in Section 2.1 to the exotic four-manifold $\mathcal{R}$ in the second item of Theorem 10.
For any integer $p\geq 1$, we define the closed oriented four-manifold
(3.1) $\mathcal{R}_{2p}=(\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}))\cup S_{p},$
@@
(3.2) $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}},$
@@
where $\alpha_{\Sigma_{2}}\subset\Sigma_{2}$ is a simple loop whose homotopy class generates the fundamental group $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2\cong\pi_{1}(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})$. The second clause of Theorem 10 implies that there is a homeomorphism
$(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})\setminus\nu(\alpha_{\Sigma_{2}})\rightarrow\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}).$
Such homeomorphism can be extended to a homeomorphism as in (3.2) by the decomposition of $\Sigma_{2p}$ in (3.3) and Corollary 5. Alternatively, the existence of the homeomorphism (3.2) also follows from work of Hambleton-Kreck [23, Theorem C].
To distinguish the diffeomorphism class of $\mathcal{R}_{2p}$ from the one of the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$, we look at the double cover
@@
that corresponds to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p=\pi_{1}(\mathcal{R}_{2p})$.
@@
If $p\in\mathbb{Z}_{>0}$ is odd, the degree-two cover of $\mathcal{R}_{2p}$ that corresponds to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to the four-manifold
@@
where $\mathcal{E}$ is the closed four-manifold of the first clause of Theorem 10, $\gamma\subset\mathcal{E}$ is a simple loop and $S_{p}$ is the compact four-manifold defined in (2.1).
[Proof]The proposition follows from the definition of $\mathcal{R}$ in [32] as the quotient of $\mathcal{E}$ by a fixed-point-free orientation-reversing involution, together with Lemma 8.
@@ 3.2. Diffeomorphism obstruction
In light of the definition of $\mathcal{E}$ in [32, Section 6.2], Proposition 12 implies the existence of a decomposition
@@
where $V=C_{0}\cup_{\phi_{0}\circ\sigma}Z$ is the building block defined in [32, Construction 6.4]. The compact four-manifold $C_{0}$ is contractible, $Z$ is obtained by attaching five 1-1-framed 22-handles to $\partial C_{0}$, the boundary $\partial V$ is the result of 0-surgery on $S^{3}$ along the Stevedore knot and $\zeta$ is any of its orientation-reversing fixed-point-free involutions. On the other hand, for every $p\geq 1$ we have that
@@
where $C_{0}^{p}=(C_{0}\setminus\nu(\gamma_{C_{0}}))\cup S_{p}$ for $\gamma_{C_{0}}\subset C_{0}$ a simple loop. We assemble
$\mathcal{E}^{\prime}_{p}=V^{\prime}\cup_{\zeta}V^{\prime}_{p},$
where $V^{\prime}=C_{0}\cup_{\phi_{0}}Z$ and $V^{\prime}_{p}=C_{0}^{p}\cup_{\phi_{0}}Z$ are made up by the same building blocks as $V$ and $V_{p}$ with different gluing maps. We refer the reader to [32, Construction 6.4] for further details.
@@
    $X_{p}=(({\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})\mathbin{\#}\Sigma_{p},$
    where $\gamma\subset{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ is a simple loop.
@@
We proceed to provide some technical lemmas that are needed to distinguish the smooth structures of Theorem C. These results are generalizations of [32, Proposition 6.9] and [32, Lemma 6.3], and make use of the Ozsváth-Szabó four-manifold invariants [35, Section 2.4]. Before delving into the lemmas and for the convenience of the reader, we briefly recall the definition of the Ozsváth–Szabó mixed invariant for closed smooth four-manifolds. The field with two elements is denoted by $\mathbb{F}$ in the sequel.
Given a closed oriented three-manifold $Y$ with a $\operatorname{Spin}^{\mathbb{C}}$-structure $\mathfrak{s}$, we denote by $HF^{-}(Y,\mathfrak{s})$ the minus version of its Heegaard Floer homology. This is a finitely generated graded module over $\mathbb{F}[U]$, whose $U$-torsion part is denoted by $HF_{\text{red}}(Y,\mathfrak{s})$.
Example 14.
The three-sphere has a unique $\operatorname{Spin}^{\mathbb{C}}$-structure $(S^{3},\mathfrak{s})$ and
$HF^{-}(S^{3},\mathfrak{s})\cong\mathbb{F}[U]$
where the degree of the unit element is $\text{deg}(1)=0$.
Given a $\operatorname{Spin}^{\mathbb{C}}$-cobordism $(W,\mathfrak{t})$ between $(Y_{1},\mathfrak{s}_{1})$ and $(Y_{2},\mathfrak{s}_{2})$, there is an induced map
$F^{-}_{(W,\mathfrak{t})}\colon HF^{-}(Y_{1},\mathfrak{s}_{1})\longrightarrow HF^{-}(Y_{2},\mathfrak{s}_{2}).$
Moreover, there is a canonical projection
$\Pi\colon HF^{-}(Y,\mathfrak{s})\cong HF_{\text{red}}(Y,\mathfrak{s})\oplus\mathbb{F}[U]^{k}\longrightarrow HF_{\text{red}}(Y,\mathfrak{s}),$
see [35, Section 2.3]. For more details on this invariant, we refer the reader to [33].
Definition 15.
Let $(W,\mathfrak{t})$ be a compact $\operatorname{Spin}^{\mathbb{C}}$-structure with non-empty boundary $\partial(W,\mathfrak{t})=(Y,\mathfrak{s})$, where $\mathfrak{s}=\mathfrak{t}|_{Y}$ is non-torsion (i.e. its first Chern class $c_{1}(\mathfrak{s})$ is not a torsion element in $H_{2}(Y;\mathbb{Z})$). The relative invariant of $(W,\mathfrak{t})$ is
$\Psi_{W,\mathfrak{t}}=\Pi(F^{-}_{(W\setminus D^{4},\mathfrak{t}|)}(1))\in HF^{-}_{\text{red}}(Y,\mathfrak{s}).$
The duality between $HF_{\text{red}}(Y,\mathfrak{s})$ and $HF_{\text{red}}(\overline{Y},\mathfrak{s})$ [36, page 376] allows us to define a pairing
$\langle\cdot,\cdot\rangle_{Y}\colon HF_{\text{red}}(Y,\mathfrak{s})\otimes HF_{\text{red}}(\overline{Y},\mathfrak{s})\longrightarrow\mathbb{F}.$
With this, we are ready to define the Ozsváth–Szabó mixed invariant for smooth four-manifolds.
Let $X$ be a closed oriented four-manifold with $b_{2}^{+}(X)=1$ and indefinite intersection form, and let $L\subseteq H_{2}(X;\mathbb{Q})$ be a one-dimensional subspace on which the intersection form vanishes. We say that a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{t})$ is allowable if $c_{1}(\mathfrak{t}|_{L})$ is non-vanishing. An embedded, oriented, separating three-manifold $Y\subseteq X$ is an admissible cut if the image in $H_{2}(X;\mathbb{Q})$ of the map induced by the inclusion $Y\rightarrow X$ is equal to $L$. At this point, we can write
$X=W_{1}\cup_{Y}W_{2}$
where $W_{1}$ and $W_{2}$ are oriented in such a way that $\partial W_{1}=Y=\overline{\partial W_{2}}$.
Definition 16.
The closed four-manifold invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with respect to the one-dimensional subspace $L\subseteq H_{2}(X;\mathbb{Q})$ is
$\Phi_{X,\mathfrak{t},L}=\langle\Psi_{(W_{1},\mathfrak{t}_{1})},\Psi_{(W_{2},\mathfrak{t}_{2})}\rangle_{Y},$
where $\mathfrak{t}_{1}=\mathfrak{t}|_{W_{1}}$, $\mathfrak{t}_{2}=\mathfrak{t}|_{W_{2}}$ and $\mathfrak{s}=\mathfrak{t}|_{Y}$.
We refer the reader to [35] and [32, Section 6.1] for further background on these invariants. We now go into the aforementioned lemmas that are needed to obstruct the existence of the diffeomorphisms of Theorem C and distinguish the smooth structures.
@@
[…]
@@
where $X_{0}(Q_{0})$ is the 0-trace of the Stevedore knot and $\gamma\subset X_{0}(Q_{0})$ is a null-homotopic simple loop.
[Proof]The case $p=1$ is [32, Proposition 6.9]. Since $C_{0}^{p}\setminus B^{4}$ is a rational homology cobordism between $S^{3}$ and $\partial C_{0}^{p}$, the same proof holds for $p>1$ by replacing $C_{0}$ with $C_{0}^{p}$ in the argument.
@@
Let $K$ be a fibered amphichiral knot of genus $g$. Let $W_{1}$ and $W_{2}$ be compact oriented 4-manifolds with positive second Betti number and boundary $\partial W_{1}=S^{3}_{0}(K)=\partial W_{2}$. Suppose there are $\operatorname{Spin}$$^{\mathbb{C}}$ structures $\{(W_{i},\mathfrak{u}_{i}):i=1,2\}$ such that $c_{1}(\mathfrak{u}_{i})$ evaluates to 2g22g-2 on a generator of $H_{2}(S^{3}_{0}(K))$ and for which $\Psi_{W_{i},\mathfrak{u}_{i}}\neq 0$. Let $X$ be the closed four-manifold that is obtained by gluing together $W_{1}$ and $W_{2}$ via any choice of orientation-reversing diffeomorphism of $S^{3}_{0}(K)$, and let $L\subset H_{2}(X;\mathbb{Q})$ be the line corresponding to the inclusion
@@
[…]
The proof of Lemma 18 is a straightforward adaptation of the one of [32, Lemma 6.3]. We now use the previous pair of lemmas to prove Proposition 13.
[Proof of Proposition 13] Item (1) follows from the construction and the fact that $\mathcal{E}^{\prime}$ is an exotic ${\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ by [32, Theorem 6.7]. By Lemma 7, there is a degree-two cover
$X_{p}\cong\mathcal{E}^{\prime}_{p}\xrightarrow{2:1}\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}.$
Let us now prove Item (2). In particular, $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ are homeomorphic by construction, since $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are homeomorphic by [32, Theorem 6.7]. To conclude the proof, we need to show that the degree-two covers $\mathcal{E}_{p}\rightarrow\mathcal{R}_{2p}$ and $\mathcal{E}^{\prime}_{p}\rightarrow\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ are not diffeomorphic. Consider the decomposition
@@
where $V_{a}$ and $V_{b}^{p}$ are copies of $V$ and $V_{p}$ respectively without the $T^{-}_{Q_{0}}$ component, see [32, Construction 6.4]. As for the case of $\mathcal{E}$, we see that $H_{2}(\mathcal{E}_{p})$ has a basis
@@
  • $(\Xi,\Theta)$ is the basis for $H_{2}(T_{Q_{0}})$ from [32, Remark 3.11]. In particular, $\Xi$ is represented by a capped-off Seifert surface for $Q_{0}$ in $S^{3}(Q_{0})$ and $\Theta$ is represented by a $2$-sphere of self-intersection 2-2 and $\Xi\cdot\Theta=1$.
[…]
$\mathcal{E}_{p}\rightarrow X_{p},$
then one could build a spherical representative for $\Xi$ as in the proof of [32, Theorem 6.7]. This would lead to a contradiction.
@@
If $p$ is even, then the four-manifold $D_{p}$ of Lemma 2 has two boundary components and the double cover $\mathcal{E}_{p}$ of $\mathcal{R}_{2p}$ is
@@
We have not been able to distinguish the smooth structure $\mathcal{R}_{2p}$ using the Heegaard Floer invariants of $\mathcal{E}_{p}$ as in [32].